OpenStudy (anonymous):

the circumference of a circle is 5 pi what is the diameter?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf C=5π }\) (this is what you are given ) \(\Large\color{blue}{ \bf C=dπ }\) (this is the formula for the circumference in terms of diameter and π ) the diameter equals to ??

OpenStudy (anonymous):

circumference=(2)(pi)(r) 5=(2)(pi)(r) 2.5=pi(r) 2.5/pi=r so when you do that, you will get the radius, when you get that do x2 to get the diameter.

OpenStudy (solomonzelman):

dg, just read what I wrote, it is as simple as that, don't make it more complicated.